OpenStudy (anonymous):

Suppose a 1,600 kg car is traveling at 20.0 m/s. What average force is needed to stop the car in 4.0 s?

OpenStudy (ash2326):

@jamalj5 We'll use the equation \[v-u=at\] here v= final velocity 0 m/s u=initial velocity 20m/s a= acceleration or deceleration in m/s^2 t= time in seconds

OpenStudy (ash2326):

Can you find a from here @jamalj5 ?

OpenStudy (anonymous):

jelly jelly penutbutter

OpenStudy (anonymous):

In my belly yeah uh yeah

OpenStudy (anonymous):

People all over the world (everybody) Join hands (join) Start a love train, love train People all over the world (all the world, now) Join hands (love ride) Start a love train (love ride), love train

OpenStudy (anonymous):

Ain't no flutterin' with the song dynasty, J. Valentine and Twista Spittin' to the rhythm while we talkin' 'bout a freak and represent the Bay area {Verse 1} Walked up in this feather, feelin' sleek-feathered one rich Got my girl dancin' like a stripper Get her all alone, hope her back is strong 'Cause I'm about to show her big dipper She look so good I wanna taste, yeah, yeah Throw it in my face, yeah, yeah I love the way she move her waist, yeah, yeah She drop it to that 808 and all that I could say is {Chorus} Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties She look so good I wanna taste, yeah, yeah Throw it in my face, yeah, yeah Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties {Verse 2} She got hella tricks, call 'em on a flip Damn that booty shakin' like Jell-o, Jell-o I'ma bone in between, call 'em super freak They jumped out on me like hello, hello She look so good I wanna taste, yeah, yeah Throw it in my face, yeah, yeah I love the way she move her waist, yeah, yeah She drop it to that 808 and all that I could say is {Chorus} Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties She look so good I wanna taste, yeah, yeah Throw it in my face, yeah, yeah Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties {Verse 3} Baby let me murder the fatty like a assassin, I wanna own it like a asset Then bend it over for some easy access, then see the booty from another aspect Girl I see your body is intricate, your beauty is infinite I wanna see her with Twista, I freak her from behind before I enter it I be so intimate 'cause I wanna suck them nipples Lookin' good girl in your Victoria Secret Now you gotta take it off, get it in like uh, uh, uh 'Cause I gotta break you off so I Gotta show her how a mack act, kill her when I pull up in the black Lacc When I get to blowin' on the fat sack, hit it from the side then the front then the back {Chorus} Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties She look so good I wanna taste, yeah, yeah Throw it in my face, yeah, yeah Shorty got ah, ah Shorty got ah, ah Shorty got retriceand titties

OpenStudy (anonymous):

what the??!? :O